OpenStudy (anonymous):

PLEASEEEEEEEEEE HELPPPPPP! Using a sheet of graph paper, solve the following system of equations graphically. Be sure to show any work, use a straight edge, and be neat. When you are finished, enter the solution set in the box below and turn in your graph paper to your teacher. x + y = 0 x - y + 2 = 0

OpenStudy (anonymous):

rewrite equation so it is in "y=" form x+y=0 [subtract x from both sides] y=-x x-y+2=0 [add y to both sides] x+2=y next, graph these two equations remember that y=mx+b b is y-intercept, that means, its where the line will cross the y axis

OpenStudy (anonymous):

x = -1 y = 1

OpenStudy (anonymous):

gohan, you should know better than to give answers, she needs the training

OpenStudy (anonymous):

First take the line x+ y = 0 & find the x intercept and y intercept . To find x - intercept put y =0 in equation x+y = 0 and to find y - intercept put x= 0, & with the help of these intercepts draw the line , similarly do for second equation x - y + 2 = 0 & draw the line .Then point of intersection of these lines is solution ....

OpenStudy (anonymous):

ok